An eight-year old boy named Nathan wants to compare his age with of his mother and grandfather. His mother is exactly four times his age and his grandfather is exactly seven timess his age. How old is Nathan's mother? How old is his grandfather? Can you rewrtie the problem statement in the form of an equation? If so, what is required for you to write it? Nathan has two brothers, Conner and Josh. If we add 2 to Conner's age, it is exactly 1/2 the age of their older brother Josh, who is 12. How old is Conner? How many years will it be until Conner's age is 1/2 of Josh's age?

Answer by richwmiller(17219)

n=8
m=4*8
g=7*8=56
c+2=j/2
j=12
c+2=6
c=4
4+x=(12+x)/2
8+2x=12+x
x=4
in four years connor will be half of josh's age
check
4+4=16/2
8=8
ok